A body of mass 12kg is projected vertically upwards and made a trajectory with base length 8m for 3secs when the body us projected vertically downwards it droped into a cylinder containing a liquid of relative density 0.92 if the distance traveled by the body vertically downwards is given as y = 3t4-4t3-8)m

Find the weight of the body

Density of the body


1
Expert's answer
2020-01-27T09:22:38-0500

Given that mass of the object = 12kg

weight of the body"= mg"

a) weight of the body "=12\times 9.8=117.6N"


Now,

Drag force due to liquid

"y=3t^4\u22124t^3\u22128"

"\dfrac{d^2y}{dx^2}=36t^2-24t"

The equation written in the last line is not clear, kindly provide the correct data, so that i can provide you the resolution.